# Black-box Membership Inference Attacks against Fine-tuned Diffusion Models

### v1

@@ Appendix B More Details for Classifier-free Guidance
As the field continues to evolve, diffusion models have the capability to generate content based on user-provided conditional prompts, a key mechanism of which is classifier-free guidance [19]. Many diffusion models, including Imagen [46], DALL·E 2 [40], and Stable Diffusion [44], which utilize the classifier-free guidance mechanism, are trained on dual objectives; however, they can be represented with a single model during training by probabilistically setting the conditional prompt to null. A conditional generation without an explicit classifier is achieved using the denoising network $\bar{\mathcal{U}}_{\theta}(x_{t},t,y)$, where
$\bar{\mathcal{U}}_{\theta}(x_{t},t,y)=(w+1)\cdot\mathcal{U}_{\theta}(x_{t},t,y)-w\cdot\mathcal{U}_{\theta}(x_{t},t).$
@@
Appendix C More Details for Traditional Black-box Attacks
In this section, we will discuss some of the most influential prior works on black-box attacks targeting image-generative models.
@@ Appendix C More Details for Traditional Black-box Attacks
Hilprecht et al. [17] argue that generative models, due to their inherent ability to capture data specifics, can exhibit overfitting and memorization. Given a query sample $x$, attackers can utilize the generative model to sample $k$ images. Define an $\epsilon$-neighborhood set $U_{\epsilon}(x)$ as $U_{\epsilon}(x)=\{x^{\prime}\mid d(x,x^{\prime})\leq\epsilon\}$. Intuitively, if a larger number of $g_{i}$ are close to $x$, the probability $\mathsf{Pr}\left[x^{\prime}\in U_{\epsilon}(x)\right]$ will also be greater. Through the Monte Carlo Integration [36], the Monte Carlo attack can be expressed as:
@@
In a further examination of the memorization and overfitting phenomena associated with generative models, Chen et al. [7] posited that the closer the generated data distribution $p_{\theta}(\hat{x})$ is to the training data distribution $q(x)$, the more likely it is for $\mathcal{M}$ to generate a query datapoint $x$. They further articulated this observation as:
$\mathsf{Pr}\left[y_{q}=1|x,\theta\right]\propto\mathsf{Pr}_{\mathcal{M}}[x|\theta_{v}]$
However, due to the inability to represent the distribution of generated data with an explicit density function, computing the precise probability becomes intractable. Consequently, Chen et al. [7] also employed the KDE method [38] and sampled $k$ times to estimate the likelihood of $x$. This can be expressed as:
$\mathsf{Pr}_{\mathcal{M}}(x|\theta)=\frac{1}{k}\sum_{i=1}^{k}K(x,\mathcal{M}(z_{i}));\quad z_{i}\sim P_{z}$ (14)
Here, $K$ denotes the kernel function, and $z_{i}$ represents the input to $\mathcal{M}$, which sample from latent code distribution $P_{z}$. Besides, Chen et al. [7] propose the approximation of Equation 14:
$\mathsf{Pr}_{\mathcal{M}}(x|\theta_{v})\approx\frac{1}{k}\sum_{i=1}^{k}\text{exp}(-d(x,\mathcal{M}(z_{i})));\quad z_{i}\sim P_{z}$ (15)
For the $k$ samples generated by $\mathcal{M}$, the distance metric $d(\cdot,\cdot)$ is employed to compute the distance between each sample $g_{i}$ and the query point $x$, and then summed up. Equation 13 and Equation 15 convey very similar meanings, for models with such undetermined sampling, the only feasible method to increase the attack success rate is through a substantial number of sampled data $g_{i}$. For a full-black attack, around $100k$ samples are required for a query point $x$ to achieve an attack AUC close to $0.60$. This undoubtedly results in significant overhead. Then, Chen et al. [7] introduced the concept of a partial-black attack. Specifically, the attacker first employs
@@
to identify the optimal latent code $z^{\ast}$. Subsequently, $g_{i}$ is generated using $\mathcal{M}(z^{\ast})$ to compare with $x$. This partial-black attack approach indeed consistently improves the success rate of attacks and reduces the number of sampled data points. However, the trade-off is the necessity to identify the optimal $z^{\ast}$. For GAN models, the input latent code $z$ is a 100-dimensional random noise. However, for contemporary conditional diffusion models, the dimensionality of the input embedding has increased exponentially. Moreover, for the latest generation of diffusion models, the input is no longer an undetermined latent code but can be explicit prompt information. Therefore, the partial-black attack tailored for GAN models is no longer applicable to these new-generation generative models.
Appendix D More Experimental Results from Varying Fine-tuning Steps
Appendix E More Experimental Results for Different Number of Inference Steps
To evaluate how inference steps affect attack accuracy, we conducted experiments on the WIT and MS COCO datasets, with results detailed in Table VI and Table VII.
TABLE VI: Influence of inference step count on MS COCO.
S Threshold-based Distribution-based Classifier-based FID
ASR AUC T@F=1% ASR AUC T@F=1% ASR AUC T@F=1%
30 $0.76$ $0.84$ $0.13$ $0.70$ $0.77$ $0.13$ $0.84$ $0.90$ $0.42$ $8.49$
50 $0.74$ $0.84$ $0.13$ $0.69$ $0.77$ $0.11$ $0.84$ $0.91$ $0.20$ $7.24$
100 $0.76$ $0.84$ $0.15$ $0.70$ $0.76$ $0.11$ $0.85$ $0.90$ $0.23$ $6.46$
200 $0.77$ $0.84$ $0.16$ $0.71$ $0.75$ $0.11$ $0.83$ $0.88$ $0.21$ $6.46$
TABLE VII: Impact of inference step count on WIT.
S Threshold-based Distribution-based Classifier-based FID
ASR AUC T@F=1% ASR AUC T@F=1% ASR AUC T@F=1%
30 $0.71$ $0.81$ $0.23$ $0.61$ $0.70$ $0.26$ $0.78$ $0.82$ $0.29$ $6.73$
50 $0.71$ $0.80$ $0.20$ $0.62$ $0.72$ $0.25$ $0.75$ $0.82$ $0.30$ $5.83$
100 $0.71$ $0.79$ $0.17$ $0.65$ $0.74$ $0.09$ $0.76$ $0.83$ $0.32$ $5.58$
200 $0.70$ $0.79$ $0.20$ $0.62$ $0.72$ $0.14$ $0.77$ $0.83$ $0.33$ $5.56$
Appendix F More Details for Comparing Five Different Image Encoders

### v4

@@ Appendix B More Details for Classifier-free Guidance
As the field advances, diffusion models can create content from user-given prompts, primarily using classifier-free guidance [19]. Many diffusion models, including Imagen [46], DALL·E 2 [40], and Stable Diffusion [44], which utilize the classifier-free guidance mechanism, are trained on dual objectives; however, they can be represented with a single model during training by probabilistically setting the conditional prompt to null. A conditional generation without an explicit classifier is achieved using the denoising network $\bar{\mathcal{U}}_{\theta}(x_{t},t,p)$, where
$\bar{\mathcal{U}}_{\theta}(x_{t},t,p)=(w+1)\cdot\mathcal{U}_{\theta}(x_{t},t,p)-w\cdot\mathcal{U}_{\theta}(x_{t},t).$
@@
Appendix C More Details for Theoretical Foundation
C-A Proof for Theorem 1
Proof:  Diffusion models employ the ELBO to approximate the log-likelihood $p(x)$ of the entire training dataset.
$\cdots$
The primary focus of optimization is on $L_{t-1}$, as explicated in the original work [18]. The other terms are treated as constants and independent decoders. The objective function can be rewritten as:
$\min\mathbf{D}_{KL}(q(x_{t-1}|x_{t},x_{0})\;\|\;p_{\theta}(x_{t-1}|x_{t})).$
Based on the assumption in DDPM [18], to elucidate further:
[…]
$\mathsf{Pr}\left[b=1|x,\theta\right]\propto-{\left\|x_{0}-\hat{x}_{\theta}(x_{t},t)\right\|^{2}_{2}}$ (10)
C-B Proof for Theorem 2
Proof:  In the original paper [44], the loss function of the Stable Diffusion model is described as follows:
The latent code $z_{t}$ is of a much smaller dimension than that of the original image. The denoising network $\mathcal{U}_{\theta}$ predicts the noise at timestep $t$ based on $z_{t}$ and the embedding generated by $\phi_{\theta}$, which takes $p$ as its input. Given that the forward process of the Stable Diffusion [44] is fixed, Equation 1 remains applicable. Therefore, by substituting in the expression $\epsilon_{t}=\frac{z_{t}-\sqrt{\bar{\alpha_{t}}}z_{0}}{\sqrt{1-\bar{\alpha}_{t}}}$ and discarding other weight terms, we can rederive the loss function of the Stable Diffusion model as:
$L_{LDM}=\mathbb{E}_{\mathcal{E}(x),t}\left[\|z_{0}-\hat{z}_{\theta}(z_{t},t,\phi_{\theta}(p))\|^{2}_{2}\right]$ (11)
As seen from Equation 11, Stable Diffusion is essentially trained to optimize image predictions at any given timestep to closely approximate the original image $\mathrm{D}(z_{0})$, where $\mathrm{D}$ is the decoder in Stable Diffusion. For the Stable Diffusion model, we can still distinguish between member samples and non-member samples by the similarity scores $\left\|\mathrm{D}(z_{0})-\mathrm{D}(\hat{z}_{\theta}(z_{t},t,\phi_{\theta}(p)))\right\|^{2}_{2}$, which is expressed as:
$\mathsf{Pr}\left[b=1|x,\theta\right]\propto-{\left\|\mathrm{D}(z_{0})-\mathrm{D}(\hat{z}_{\theta}(z_{t},t,\phi_{\theta}(p)))\right\|^{2}_{2}}$ (12)
Appendix D More Details for Traditional Black-box Attacks
@@
Hilprecht et al. [17] argue that generative models can overfit and memorize data due to their ability to capture specific details. Given a query sample $x$, attackers can utilize the generative model to sample $k$ images. Define an $\epsilon$-neighborhood set $U_{\epsilon}(x)$ as $U_{\epsilon}(x)=\{x^{\prime}\mid d(x,x^{\prime})\leq\epsilon\}$. Intuitively, if a larger number of $g_{i}$ are close to $x$, the probability $\mathsf{Pr}\left[x^{\prime}\in U_{\epsilon}(x)\right]$ will also be greater. Through the Monte Carlo Integration [36], the Monte Carlo attack can be expressed as:
@@
In a further examination of the memorization with generative models, Chen et al. [7] posited that the closer the generated data distribution $p_{\theta}(\hat{x})$ is to the training data distribution $q(x)$, the more likely it is for $\mathcal{G}$ to generate a query datapoint $x$. They further articulated this observation as:
$\mathsf{Pr}\left[y_{q}=1|x,\theta\right]\propto\mathsf{Pr}_{\mathcal{G}}[x|\theta_{v}]$
However, due to the inability to represent generated data distribution with an explicit density function, computing the precise probability becomes intractable. Therefore, Chen et al. [7] also employed the KDE method [38] and sampled $k$ times to estimate the likelihood of $x$. This can be expressed as:
$\mathsf{Pr}_{\mathcal{G}}(x|\theta)=\frac{1}{k}\sum_{i=1}^{k}K(x,\mathcal{G}(z_{i}));\quad z_{i}\sim P_{z}$ (14)
Here, $K$ denotes the kernel function, and $z_{i}$ represents the input to $\mathcal{G}$, which sample from latent code distribution $P_{z}$. Besides, Chen et al. [7] propose the approximation of Equation 14:
$\mathsf{Pr}_{\mathcal{G}}(x|\theta_{v})\approx\frac{1}{k}\sum_{i=1}^{k}\text{exp}(-d(x,\mathcal{G}(z_{i})));\quad z_{i}\sim P_{z}$ (15)
For the $k$ samples from $\mathcal{G}$, we use the distance metric $d(\cdot,\cdot)$ to measure and sum the distances between each sample $g_{i}$ and the query point $x$Equation 13 and Equation 15 indicate that the only practical way to improve attack success rates in models with such stochastic sampling is by significantly increasing the number of samples $g_{i}$. For a full-black attack, around $100k$ samples are required for a query point $x$ to achieve an attack AUC close to $0.60$. This undoubtedly results in significant overhead. Then, Chen et al. [7] introduced the concept of a partial-black attack. Specifically, the attacker first employs
@@
to identify the optimal latent code $z^{\ast}$. Subsequently, $g_{i}$ is generated using $\mathcal{G}(z^{\ast})$ to compare with $x$. The partial-black attack method boosts the success rate of attacks and uses fewer data samples, but requires finding the optimal $z^{\ast}$. In GANs, the input latent code $z$ is typically a 100-dimensional random noise. However, in newer conditional diffusion models, the complexity and size of input embeddings have greatly increased. Additionally, the latest diffusion models use explicit prompt information instead of random latent codes, making the partial-black attack strategy used for GANs unsuitable for these new-generation generative models.
Appendix E More Details for Comparing Five Different Image Encoders
@@
Appendix F More Experimental Results for Varying Fine-tuning Steps
Appendix G More Experimental Results for Different Number of Inference Steps
To evaluate how inference steps affect attack performance, we conducted experiments on the WIT and MS COCO datasets, with results detailed in Table XI.
TABLE XI: Experiment results for more inference steps on MS COCO and WIT.
S MS COCO WIT
Threshold-based Distribution-based Classifier-based FID Threshold-based Distribution-based Classifier-based FID
ASR AUC T@F=1% ASR AUC T@F=1% ASR AUC T@F=1% ASR AUC T@F=1% ASR AUC T@F=1% ASR AUC T@F=1%
$30$ $0.76$ $0.84$ $0.13$ $0.70$ $0.77$ $0.13$ $0.84$ $0.90$ $0.42$ $8.49$ $0.71$ $0.81$ $0.23$ $0.61$ $0.70$ $0.26$ $0.78$ $0.82$ $0.29$ $6.73$
$50$ $0.74$ $0.84$ $0.13$ $0.69$ $0.77$ $0.11$ $0.84$ $0.91$ $0.20$ $7.24$ $0.71$ $0.80$ $0.20$ $0.62$ $0.72$ $0.25$ $0.75$ $0.82$ $0.30$ $5.83$
$100$ $0.76$ $0.84$ $0.15$ $0.70$ $0.76$ $0.11$ $0.85$ $0.90$ $0.23$ $6.46$ $0.71$ $0.79$ $0.17$ $0.65$ $0.74$ $0.09$ $0.76$ $0.83$ $0.32$ $5.58$
$200$ $0.77$ $0.84$ $0.16$ $0.71$ $0.75$ $0.11$ $0.83$ $0.88$ $0.21$ $6.46$ $0.70$ $0.79$ $0.20$ $0.62$ $0.72$ $0.14$ $0.77$ $0.83$ $0.33$ $5.56$